three resistors of 2, 3, and 6 ohm are connected in (i)series (ii)parallel.calculate the ratio of effective resistance of series and parallel combination of resistances
1
Expert's answer
2012-07-06T07:31:03-0400
(i) resistors are connected in series:

Rs = 2 + 3 + 6 = 11 ohm.

(ii) resistors are connected in parallel:

Rp = 1/(1/2 + 1/3 + 1/6) = 1 ohm.

The ratio is Rs/Rp = 11/1 = 11.